OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (shane_b):

Do you have an idea on where to start?

OpenStudy (anonymous):

nope not at all lol

OpenStudy (shane_b):

I'd start with the left part of the equation. Get rid of the parentheses by distributing the -2. Then do the same for the right side. After that, it will become pretty simple to move things around and solve for x.

OpenStudy (anonymous):

so u multiply - 2*x= -2 and -2*3= -6

OpenStudy (auctoratrox):

so then you get -2x - 6 on the left by using the distributive property. on the right you get -2x - 2 - 4 because of the distributive property. combine like terms and you get 0 = 0, which is a true statement! therefore x can take on infinitely many values and still satisfy the equation.

OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 -2x-6 = -2x-2-4 -2x+2x= -6 +6 0=0 it is a true statement!